# Tony Lindeberg

### v1

@@ 5.1.4 The special case with equal image velocities
An explanation for this simplicity of the resulting relationships is because of the Galilean covariance property of spatio-temporal receptive field model, which then up to a Galilean transformation is isomorphic to a purely space-time separable model for the spatio-temporal receptive fields based on spatio-temporal smoothing operations of the form
$\begin{split}T(x,t;\;s,\Sigma,\tau=g(x;\;s,\Sigma)\,h(t;\;\tau)=\end{split}$
@@
In the treatment so far, we have made use of special algebraic properties as arising from using spatial and spatio-temporal smoothing kernels solely based on Gaussian kernels, for deriving both infinitesimal and macroscopic relationships between receptive field responses computed for different values of the filter parameters.
When replacing the non-causal temporal Gaussian kernel $h(t;\;\tau)=g(t;\;\tau)$ according to (3) by the time-causal limit kernel $h(t;\;\tau)=\psi(t;\;\tau,c)$ according to (4), where $c>1$ is a distribution parameter that specifies the ratio between the temporal scale level in units of the temporal standard deviation of the kernel, we obtain a joint spatio-temporal smoothing kernel of the form
@@
From these structural differences arising from using the time-causal limit kernel as the temporal smoothing kernel instead of the non-causal temporal Gaussian kernel, we cannot expect to be able to directly apply similar methodologies for deriving infinitesimal or macroscopic evolution properties over the filter parameters for the time-causal spatio-temporal kernel (171) as we used for deriving such relationships for the fully Gaussian-smoothing based receptive field models in Sections 4 and 5.
@@
Figure 9: Generalized Hermite polynomials as arising from derivatives of the time-causal spatio-temporal kernel $T(x,t;\;s,\Sigma,\tau,v,c)$ according to (171), based on smoothing with a spatially anisotropic affine Gaussian kernel, with respect to only the image coordinates $x=(x_{1},x_{2})^{T}$ up to order 2, as well as with respect to only the spatial scale parameter $s$, the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$.
Regarding infinitesimal relationships, the differential relationships that we derived in Section 4.4 regarding derivative operators solely expressed over the spatial domain will also hold if we replace the non-causal Gaussian kernel $h(t;\;\tau)=g_{1D}(t;\;\tau)$ in (119) by the time-causal limit kernel $h(t;\;\tau)=\psi(t;\;\tau,c)$. Thus, if we for any spatio-temporal video stream $f\colon{\mathbb{R}}^{r}\times{\mathbb{R}}\rightarrow R$ define a time-causal spatio-temporal scale-space representation according to
This can be easily understood by observing that when calculating the ratios between derivatives with respect to either the spatial coordinates $x$ or the spatially based filter parameters $s,\Sigma)$, the occurence of the time-causal limit kernel will cancel in the resulting expressions. Thus, we obtain similar results for these ratios computed from the time-causal spatio-temporal smoothing kernel as shown in Figure 9, as we previously obtained for the corresponding non-causal spatio-temporal smoothing kernel in Figure 8.
$L(\cdot,\cdot;\;s_{j},\Sigma_{j},\tau_{j},v_{j},c)=\\ =\Delta T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\tau_{i},\Delta v,c)*L(\cdot,\cdot;\;s_{i},\Sigma_{i},\tau_{i},v_{i},c),$ (191)
where the incremental kernel $\Delta T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\tau_{i},\Delta v,c)$ is given by
$\begin{split}\Delta T(x,t;\;\Delta s,\Sigma,\tau_{i},\Delta v,c)=\end{split}$
$\begin{split}=g(x-\Delta v\,t;\;\Delta s,\Delta\Sigma)\,h_{\exp}(t;\;\sqrt{c^{2}-1}\sqrt{\tau_{i}})\end{split}$
$\begin{split}\hphantom{=}\,\,h_{\exp}(t;\;\sqrt{c^{2}-1}\sqrt{\tau_{i}})\end{split}$ (192)
here, for simplicity, restricted to adjacent temporal scales
$\tau_{i}=\frac{\tau_{j}}{c^{2}}.$ (193)
The motivation for the last relationship is that the temporal scale levels in these types of temporal and spatio-temporal scale-space representations are inherently discrete according to a geometric distribution of the form (6).
6.2.1 The special case with equal image velocities
Unfortunately, it appears as a rather complex problem to derive closed-form expressions for the relationships between the parameters $(\Delta s,\Delta\Sigma,\tau_{i},\Delta v)$ of the incremental kernel from the parameters $(s_{i},\Sigma_{i},\tau_{i},v_{i})$ and $(s_{j},\Sigma_{j},\tau_{j},v_{j})$ of two involved layers of receptive field responses in the general case, when the image velocities $\Delta v$, $v_{i}$ and $v_{j}$ may be different. A main reason to technical complications is that the time-causal limit kernel does not have any compact explicit expression, since it is given as the convolution of an infinite number of truncated exponential kernels.
In the special case when the image velocities are all equal
$v_{j}=\Delta v=v_{i},$ (194)
$\tau_{i}=\frac{\tau_{j}}{c^{2}},$ (195)
7 Summary and discussion
After an overview of the generalized Gaussian derivative model for visual receptive fields in Section 3, with emphasis on its relation to handle variabilities in image structures as generated by natural geometric image transformations as modelled by local linearizations in terms of (i) uniform spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, we have specifically, from the requirement of covariance properties of the receptive fields under such geometric image transformations, motivated the importance of multi-parameter families of visual receptive fields. This motivation originates from the desirable requirement of making it possible to up to first order of approximation match the receptive field responses that have been computed from image data acquired under different viewing conditions. The requirement of closedness of the representation of receptive field responses under such geometric image transformations then means that the shapes of the receptive fields ought to be expanded over the degrees of freedom of the corresponding geometric image transformations.
For our considered axiomatically determined normative theory of visual receptive fields, this leads to spatial and spatio-temporal receptive fields, that are based on applying spatial and/or spatio-temporal derivative operators to a multi-parameter spatial and/or spatio-temporal scale-space representation of the either purely spatial or joint spatio-temporal image data. Specifically, we have, by the covariance properties of the receptive fields in Equations (29)–(30) and (38)–(39) related the variabilities in the shapes of the receptive fields to the parameters of the general class of composed geometric image transformations of the form (33) and (34), which as illustrated in Figure 4 represents the first-order variability in image structures when observing the surfaces of smooth objects in dynamic scenes.
Specifically, in Section 4, we studied three variants of models for spatial or spatio-temporal smoothing operations in the receptive fields based on pure Gaussian smoothing over the spatial or spatial-temporal domains, in terms of either (i) a purely spatial model with 3 effective degrees of freedom in the parameter space, (ii) a joint spatio-temporal model with a restriction to isotropic Gaussian smoothing over the spatial domain, resulting in 4 effective degrees of freedom in parameter space, and (ii) a joint spatio-temporal model with general affine Gaussian smoothing over the spatial domain, leading to 6 effective degrees of freedom in the parameter space. By the special algebraic structures of the resulting receptive field representations, when using solely Gaussian smoothing over both the spatial and the temporal domains, we could introduce the notion of generalized Hermite polynomials of the resulting spatial or spatio-temporal smoothing operations to substantially simplify the process of deriving and proving the corresponding differential evolution equations over infinitesimal perturbations of the filter parameters, largely corresponding to the notion of infinitesimal generators for more traditional semi-groups.
By these derived evolution properties over variations of the filter parameters of the receptive fields, we do foremost obtain a better theoretical understanding about how receptive field responses are related between different values of the filter parameters. These either infinitesimal or macroscopic evolution properties can, in turn, be used in computational schemes to compute banks of receptive field responses suitable for matching the responses of receptive fields applied to image data that have been acquired under substantial variations in the viewing conditions, which according to the theory described in Section 3.5 implies that the shapes of the receptive fields would have to be expanded over the degrees of freedom of the corresponding geometric image transformations. Thereby, it becomes possible to, to first order of approximation, match the receptive field responses between different views of the same object or a similar spatio-temporal event, and thus establish an identity operation between the receptive field responses computed for different observations of the same object or a similar spatio-temporal event.
These theoretical results do also have possible implications for theories about biological vision, where indeed such a model concerning an expansion over receptive field shapes over the degrees of freedom corresponding to the degrees of freedom in the here studied generalized Gaussian derivative model for visual receptive fields has in Lindeberg (2025c) been recently proposed for predicting and explaining variabilities of simple cells in the primary visual cortex of higher mammals. From statistics of the populations of neurons in the early visual pathway, with about 1 M output channels from the retina are mapped to 1 M output channels from the lateral geniculate nucleus (LGN) to the primary visual cortex (V1), to about 190 M neurons in V1 with 37 M output channels (see DiCarlo et al. (2012) Figure 3), such a substantial expansion of the number of receptive fields from the LGN to V1 would indeed be consistent with an expansion of the shapes of the receptive fields over shape parameters of the receptive fields.
Specifically, the evolution properties between receptive field responses for different sets of filter parameters derived in this paper provide a theoretical explanation for the alternative design strategy of a representation of the receptive field responses of simple cells proposed in Lindeberg (2025c), where an idealized vision system could by the use of cascade smoothing properties over the spatial and temporal scale parameters choose to only implement receptive field responses corresponding to the finest spatial and temporal scale levels in the sensorium, corresponding to the primary visual cortex of higher mammals. Then, at higher levels in the visual hierarchy, equivalent receptive field responses corresponding to coarser levels of spatial and temporal scales could be computed in an indirect manner, either based on integration of the infinitesimal evolution equations derived in Section 4 or based on the macroscopic evolution properties derived in Section 5.

### v2

@@ 5.1.4 The special case with equal image velocities
An explanation for this simplicity of the resulting relationships is because of the Galilean covariance property of the spatio-temporal receptive field model, which then up to a Galilean transformation is isomorphic to a purely space-time separable model for the spatio-temporal receptive fields based on spatio-temporal smoothing operations of the form
$\begin{split}T(x,t;\;s,\Sigma,\tau)=g(x;\;s,\Sigma)\,h(t;\;\tau)=\end{split}$
@@
In the treatment so far, we have made use of special algebraic properties, as arising from using spatial and spatio-temporal smoothing kernels solely based on Gaussian kernels, for deriving both infinitesimal and macroscopic relationships between the receptive field responses computed for different values of the filter parameters.
When replacing the non-causal temporal Gaussian kernel $h(t;\;\tau)=g(t;\;\tau)$ according to (3) by the time-causal limit kernel $h(t;\;\tau)=\psi(t;\;\tau,c)$ according to (4), where $c>1$ is a distribution parameter that specifies the ratio between the temporal scale levels in units of the temporal standard deviation of the kernel, we obtain a joint spatio-temporal smoothing kernel of the form
@@
From these structural differences, arising from using the time-causal limit kernel as the temporal smoothing kernel instead of the non-causal temporal Gaussian kernel, we cannot expect to be able to directly apply similar methodologies for deriving infinitesimal or macroscopic evolution properties over the filter parameters for the time-causal spatio-temporal kernel (138), as we used for deriving such relationships for the fully Gaussian-smoothing based receptive field models in Sections 4 and 5.
@@
Regarding infinitesimal relationships, the differential relationships, that we derived in Section 4.4 regarding derivative operators solely expressed over the spatial domain, will also hold if we replace the non-causal Gaussian kernel $h(t;\;\tau)=g_{1D}(t;\;\tau)$ in (86) by the time-causal limit kernel $h(t;\;\tau)=\psi(t;\;\tau,c)$. Thus, if we for any spatio-temporal video stream $f\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\rightarrow{\mathbb{R}}$ define a time-causal spatio-temporal scale-space representation according to
This can be easily understood by observing that when calculating the ratios between derivatives with respect to either the spatial coordinates $x$ or the spatially based filter parameters $(s,\Sigma)$, the occurence of the time-causal limit kernel will cancel in the resulting expressions. Thus, we obtain similar results for these ratios computed from the time-causal spatio-temporal smoothing kernel as shown in Figure 9 in Appendix C, as we previously obtained for the corresponding non-causal spatio-temporal smoothing kernel shown in Figure 8 in Appendix B.
$L(\cdot,\cdot;\;s_{j},\Sigma_{j},\tau_{j},v_{j},c)=\\ =\Delta T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\Delta\tau,\Delta v,c)*L(\cdot,\cdot;\;s_{i},\Sigma_{i},\tau_{i},v_{i},c),$ (148)
where the incremental kernel $\Delta T(\cdot,\cdot;\;\Delta s,\Delta\Sigma,\Delta\tau,\Delta v,c)$ is given by
$\begin{split}\Delta T(x,t;\;\Delta s,\Sigma,\Delta\tau,\Delta v,c)=\end{split}$
$\begin{split}=g(x-\Delta v\,t;\;\Delta s,\Delta\Sigma)\,h_{\exp}(t;\;\sqrt{\Delta\tau})\end{split}$
$\begin{split}\hphantom{=}\,\,h_{\exp}(t;\;\sqrt{\Delta\tau})\end{split}$ (149)
here, for simplicity, restricted to adjacent temporal scales
$\tau_{i}=\frac{\tau_{j}}{c^{2}}$ (150)
with
$\Delta\tau=\tau_{j}-\tau_{i}=\frac{c^{2}-1}{c^{2}}\,\tau_{j}.$ (151)
The motivation for the last relationships is that the temporal scale levels in these types of temporal and spatio-temporal scale-space representations are inherently discrete according to a geometric distribution of the form (6).
6.2.1 The special case with equal image velocities
Unfortunately, it appears as a rather complex problem to derive closed-form expressions for the relationships between the parameters $(\Delta s,\Delta\Sigma,\Delta\tau,\Delta v)$ of the incremental kernel from the parameters $(s_{i},\Sigma_{i},\tau_{i},v_{i})$ and $(s_{j},\Sigma_{j},\tau_{j},v_{j})$ of two involved layers of receptive field responses in the general case, when the image velocities $\Delta v$, $v_{i}$ and $v_{j}$ may be different. A main reason to technical complications is that the time-causal limit kernel does not have any compact explicit expression, since it is given as the convolution of an infinite number of truncated exponential kernels.
In the special case when the image velocities are all equal
$v_{j}=\Delta v=v_{i},$ (152)
$\tau_{i}=\frac{\tau_{j}}{c^{2}},$ (153)
7 Summary and discussion
After an overview of the generalized Gaussian derivative model for visual receptive fields in Section 3, with emphasis on its relation to handle variabilities in image structures, as generated by natural geometric image transformations, and as modelled by local linearizations in terms of (i) uniform spatial scaling transformations, (ii) spatial affine transformations, (iii) Galilean transformations and (iv) temporal scaling transformations, we have specifically, from the requirement of covariance properties of the receptive fields under such geometric image transformations, motivated the importance of multi-parameter families of visual receptive fields. This motivation originates from the desirable requirement of making it possible to, up to first order of approximation, match the receptive field responses that have been computed from image data acquired under different viewing conditions. The requirement of closedness of the representation of receptive field responses under such geometric image transformations then means that the shapes of the receptive fields ought to be expanded over the degrees of freedom of the corresponding geometric image transformations.
For our considered axiomatically determined normative theory of visual receptive fields, this leads to spatial and spatio-temporal receptive fields, that are based on applying spatial and/or spatio-temporal derivative operators to a multi-parameter spatial and/or spatio-temporal scale-space representation of the either purely spatial or joint spatio-temporal image data. Specifically, we have, by the covariance properties of the receptive fields in Equations (30)–(31) and (39)–(40) related the variabilities in the shapes of the receptive fields to the parameters of the general class of composed geometric image transformations of the form (34) and (35), which, as illustrated in Figure 4, represents the first-order variability in image structures, when observing the surfaces of smooth objects in dynamic scenes.
Specifically, in Section 4, we studied three variants of models for spatial or spatio-temporal smoothing operations in the receptive fields based on pure Gaussian smoothing over the spatial or spatial-temporal domains, in terms of either (i) a purely spatial model with 3 effective degrees of freedom in the parameter space, (ii) a joint spatio-temporal model with a restriction to isotropic Gaussian smoothing over the spatial domain, resulting in 4 effective degrees of freedom in the parameter space, and (ii) a joint spatio-temporal model with general affine Gaussian smoothing over the spatial domain, leading to 6 effective degrees of freedom in the parameter space. By the special algebraic structures of the resulting receptive field representations, when using solely Gaussian smoothing over both the spatial and the temporal domains, we could introduce the notion of generalized Hermite polynomials of the resulting spatial or spatio-temporal smoothing operations, to substantially simplify the process of deriving and proving the corresponding differential evolution equations over infinitesimal perturbations of the filter parameters, largely corresponding to the notion of infinitesimal generators for more traditional semi-groups.
From the viewpoint of the underlying normative theory for visual receptive fields, the receptive field shapes are parameterized by over the 4-D (although effectively 3-D) parameters $(s,\Sigma)$ in the purely spatial case and by the 7-D (although effectively 6-D) parameters $(s,\Sigma,\tau,v)$ in the joint spatio-temporal case. If we regard this theory as an idealized model of the visual processing in the primary visual cortex, then flattening out these higher-dimensional spaces onto the 2-D cortical surface, implies that points $p$ and parameters $P$ that may be relatively nearby in the higher-dimensional spaces $(p;\;P)=(x;\;s,\Sigma)$ or $(p;\;P)=(x;\;s,\Sigma,\tau,v)$ may be further away from each other when flattened out on the 2-D cortical surface.
In this way, if we transfer the cascade smoothing properties derived in this paper onto such a flattened cortical manifold, then the corresponding connections between receptive fields at different positions on the cortical surface will lead to a set of “horizontal connections” between different neurons. Specifically, due to the flattening of the higher-dimensional manifold of parameterized receptive responses onto the 2-D cortical surface, some of these connections will by necessity be “long range” on the cortical surface, although they would have been of shorter range in the original higher-dimensional manifold of receptive field responses. In this respect, the presented theory offers a possible explanation of a certain kind of “horizontal connections” constructed from a different purpose than the more commonly used explanation of integrating receptive field responses at different positions in the visual field into higher-order “gestalt” entities, such as contours (see e.g. Bolz and Gilbert (1989), Bosking et al. (1997), Ben-Shahar and Zucker (2004) and Sarti et al. (2009)). [^6] Let us, however, remark that with this statement we do not in any way intend to argue against the use of “horizontal connections” between different visual neurons for computing higher-order “gestalt” entities, such as contours. Instead, our purpose is to propose a complementary explanation regarding a possible subset of such “horizontal connections”.
For modelling visual receptive fields, there is, as previously described in Section 2, also an alternative school of using idealized models based on Gabor functions, which consist of the multiplication of Gaussian functions with complex sine waves. For the purely Gaussian-smoothing-based component in the Gabor model, it seems likely that structurally related evolution properties over the receptive field parameters could be obtained for the corresponding Gabor models of visual receptive fields. Regarding variations in the wavelength or the orientation of the complex sine wave component in the Gabor model, the corresponding relationships could, however, be expected to be of a different nature.
In this treatment, we have therefore focused solely on the topic of deriving evolution properties over the parameters of the generalized Gaussian derivative model, which obeys very special closedness properties, that could not be expected to hold for more generic classes of kernels. Thus, we leave the topic of deriving corresponding infinitesimal, or, if possible, macroscopic relationships for the Gabor model to future work. For a complementary in-depth analysis of similarities vs. differences between the generalized Gaussian derivative and the Gabor models regarding orientation selectivity properties, see Lindeberg (2025b).
By the derived evolution properties over variations of the filter parameters of the receptive fields according to the generalized Gaussian derivative model for visual receptive fields, we do foremost obtain a better theoretical understanding about how receptive field responses are related between different values of the filter parameters. These either infinitesimal or macroscopic evolution properties can, in turn, be used in computational schemes to compute banks of receptive field responses suitable for matching the responses of receptive fields applied to image data that have been acquired under substantial variations in the viewing conditions, which according to the theory described in Section 3.5 implies that the shapes of the receptive fields would have to be expanded over the degrees of freedom of the corresponding geometric image transformations. Thereby, it becomes possible to, to first order of approximation, match the receptive field responses between different views of the same object or a similar spatio-temporal event, and thus establish an identity operation between the receptive field responses computed for different observations of the same object or a similar spatio-temporal event.
These theoretical results do also have possible implications for theories about biological vision, where indeed such a model concerning an expansion over receptive field shapes over the degrees of freedom, corresponding to the degrees of freedom in the here studied generalized Gaussian derivative model for visual receptive fields, has in Lindeberg (2025d) been recently proposed for predicting and explaining variabilities of simple cells in the primary visual cortex of higher mammals. From statistics of the populations of neurons in the early visual pathway, with about 1 M output channels from the retina are mapped to 1 M output channels from the lateral geniculate nucleus (LGN) to the primary visual cortex (V1), to about 190 M neurons in V1 with 37 M output channels (see DiCarlo et al. (2012) Figure 3), such a substantial expansion of the number of receptive fields from the LGN to V1 would indeed be consistent with an expansion of the shapes of the receptive fields over shape parameters of the receptive fields.
Specifically, the evolution properties between receptive field responses for different sets of filter parameters derived in this paper provide a theoretical explanation for the alternative design strategy of a representation of the receptive field responses of simple cells proposed in Lindeberg (2025d), where an idealized vision system could by the use of cascade smoothing properties over the spatial and temporal scale parameters choose to only implement receptive field responses corresponding to the finest spatial and temporal scale levels in the sensorium, corresponding to the simple cells in the primary visual cortex of higher mammals. Then, at higher levels in the visual hierarchy, equivalent receptive field responses, corresponding to coarser levels of spatial and temporal scales, could be computed in an indirect manner, either based on integration of the infinitesimal evolution equations derived in Section 4, or based on the macroscopic evolution properties derived in Section 5.
Appendix
Appendix A Generalized Hermite polynomials for the isotropic spatio-temporal Gaussian kernel $T(x,t;\;s,\tau,v)$
Figure 7: Generalized Hermite polynomials as arising from derivatives of the isotropic spatio-temporal Gaussian kernel $T(x,t;\;s,\tau,v)$ according to (79), based on smoothing with a spatially isotropic Gaussian kernel, with respect to the image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$.
Figure 7 lists generalized Hermite polynomials for the isotropic spatio-temporal Gaussian kernel $T(x,t;\;s,\tau,v)$ according to (79).
[…]
Figure 8: Generalized Hermite polynomials as arising from derivatives of the affine spatio-temporal Gaussian kernel $T(x,t;\;s,\Sigma,\tau,v)$ according to (86), based on smoothing with a spatially anisotropic affine Gaussian kernel, with respect to the image coordinates $x=(x_{1},x_{2})^{T}$ and the time variable $t$ up to order 2, as well as with respect to the spatial scale parameter $s$, the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$, the temporal scale parameter $\tau$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$. (The explicit expressions for $\partial_{x_{1}t}T(x,t;\;s,\Sigma,\tau,v)$, $\partial_{x_{2}t}T(x,t;\;s,\Sigma,\tau,v)$ and $\partial_{tt}T(x,t;\;s,\Sigma,\tau,v)$ have been omitted here, because they do not fit within two lines each.)
Appendix B Generalized Hermite polynomials for the affine spatio-temporal Gaussian kernel $T(x,t;\;s,\Sigma,\tau,v)$
Figure 8 lists generalized Hermite polynomials for the affine spatio-temporal Gaussian kernel $T(x,t;\;s,\Sigma,\tau,v)$ according to (86).
@@ Appendix A Generalized Hermite polynomials for the isotropic spatio-temporal Gaussian kernel $T(x,t;\;s,\tau,v)$
Figure 9: Generalized Hermite polynomials as arising from derivatives of the time-causal affine spatio-temporal kernel $T(x,t;\;s,\Sigma,\tau,v,c)$ according to (138), based on smoothing with a spatially anisotropic affine Gaussian kernel, with respect to only the image coordinates $x=(x_{1},x_{2})^{T}$ up to order 2, as well as with respect to only the spatial scale parameter $s$, the elements $\Sigma_{11}$, $\Sigma_{12}$ and $\Sigma_{22}$ of the spatial covariance matrix $\Sigma$ and the elements $v_{1}$ and $v_{2}$ of the velocity vector $v$.
Appendix C Generalized Hermite polynomials for the time-causal affine spatio-temporal kernel $T(x,t;\;s,\Sigma,\tau,v,c)$
Figure 9 lists generalized Hermite polynomials for the time-causal affine spatio-temporal Gaussian kernel $T(x,t;\;s,\Sigma,\tau,v,c)$ according to (138).